please help me solve this question
Find a solution to the following system of equations.
-2x+5y=8
2x+2y=6

Answer by sachi(548)

now add both the eqns
-2x+5y=8.....1
2x+2y=6 .....2
-----------------
+7y=14
or y=14/7=2
from eqn 1
-2x+5*2=8
or -2x=8-10=-2
or x=-2/-2=1